Use logarithmic differentiation to find the derivative of the function.
y = sqrt(x)e^x^2(x^2 + 3)^5
1
Expert's answer
2011-02-28T05:27:31-0500
y' = 1/(2√x) * ex^2 (x2+3)5 + 2x√x ex^2 (x2+3)5 +2x*5* √x ex^2 (x2+3)4
= ex^2(x2+3)4 (1/(2√x)(x2+3) + 2x3√x + 6x√x + 10x√x) = ex^2(x2+3)4(1/2 * x√x +6x√x + 10x√x + 3/(2√x) + 2x3√x) =
= ex^2(x2+3)4 (16.5x√x + 3/(2√x) + 2x3√x)